OpenStudy (anonymous):

How would I convert this quadratic function to intercept and vertex form? y=4x^2-12x+5

OpenStudy (amistre64):

by completeing the square

OpenStudy (amistre64):

y = 4x^2 -12x +5 ; regroup y = (4x^2 -12x) +5 ; factor out a 4 y = 4(x^2 -3x) +5 ; and complete the square .... this is where knowing the technique is very useful

OpenStudy (amistre64):

y = 4(x^2 -3x+N) +5 -4N ; where N is the value that "completes the square" in this case its 9/4 y = 4(x^2 -3x+9/4) +5 -9 ; compact it y = 4(x-3/2)^2 - 4 and its done

OpenStudy (amistre64):

given the form: x^2 +bx; a complete square is obtained by taking half of "b"; b/2 ... and squareing it: (b/2)^2 = b^2/4

OpenStudy (amistre64):

the other trick is in once you add it, you have to subtract it back out in order to keep the same value of the original equation

OpenStudy (amistre64):

x^2 + 3x; (3/2)^2 = 9/4 ... which is what i meant

OpenStudy (anonymous):

Oh, okay...I think I get it a little better now. I always get stuck on completing the square and go downhill from there. Thank you...!

OpenStudy (amistre64):

youre welcome :)